Calculate the mass of potassium hydroxide, KOH, required to prepare a 200 mL

solution that was diluted to give a concentration of 0.030 mol/L. The solution has

a dilution factor of 5.


1
Expert's answer
2021-08-11T05:22:59-0400

V1 = 200 mL

C2 = 0.030 mol/L

Dilution factor = 5


Solution:

C1V1 = C2V2

Dilution factor = C1/C2 = V2/V1

Dilution factor = Final solution volume (V2) / Volume from stock solution (V1)

Therefore,

Final solution volume (V2) = Dilution factor × Volume from stock solution (V1)

V2 = 5 × (200 mL) = 1000 mL = 1 L


Calculate the moles of KOH:

n(KOH) = C2 × V2 = (0.030 mol/L) × (1 L) = 0.030 mol KOH


The molar mass of KOH is 56.1 g/mol

Therefore,

(0.030 mol KOH) × (56.1 g KOH / 1 mol KOH) = 1.683 g KOH = 1.68 g KOH


Answer: 1.68 g of KOH is required to prepare a 200 mL solution.